# Singular Solutions to the Complex Monge-Ampère equation

### v1

@@ Abstract.
We present an explicit pluripotential (as well as viscosity) solution to the complex Monge-Ampère equation with constant right-hand side on $\mathbb{D}\times\mathbb{C}^{n-1}\,(n\geq 2)$, which lies merely in $W^{1,2}_{loc}\cap W^{2,1}_{loc}$ and is not even Dini continuous. Additionally, we exhibit two families of explicit entire toric solutions on $\mathbb{C}^{n}$ with continuous Hölder exponent $\alpha\in(0,1)$ and $W^{2,p}_{loc}$ exponent $p\in(1,2)$.
@@
Jiaxiang Wang is partially supported by the Fundamental Research Funds for the Central Universities Nankai University (050-63241427). Wenlong Wang is partially supported by NSFC 12071338, 12001292 and the Fundamental Research Funds for the Central Universities Nankai University (050-63243065).
@@ 1. Introduction
Complex Monge-Ampère equations plays an important role in complex analysis and geometry. There are two fundamental methods for generalizing the notion of solution to functions with low regularity. One is the pluripotential approach founded by Bedford and Taylor [2], while the other is the viscosity approach introduced by Crandall and Lions for nonlinear PDEs [13] (see the well-known survey [12]) and developed for the complex Monge-Ampère equation in [15, 18, 19, 20, 28].
@@
where ¯$\partial\bar{\partial}u$ denotes the complex Hessian of $u$.
@@
Clearly, Hölder continuity implies Dini continuity. Therefore, Theorem 1.1 indicates that pure interior Hölder estimate fails for the complex Monge-Ampère equation. This theorem also shows that the equation provides almost no regularity improvement for the gradient of the solution, as a bounded plurisubharmonic function naturally lies in $W^{1,2}_{loc}$. For the Hessian, it remains unclear whether $L^{1}_{loc}$ is the lowest possible regularity, as is the case for the real Monge-Ampère equation (cf. [14, 25, 26]).
@@
Besides Example 1.2, we also construct examples that are toric and global on ${\mathbb{C}}^{n}$, i.e., functions satifying $u(z_{1},\cdots,z_{n})=u(e^{{\sqrt{-1}}t_{1}}z_{1},\cdots,e^{{\sqrt{-1}}t_{n}}z_{n})$ for any $(z_{1},\cdots,z_{n})\in\mathbb{C}^{n}$ and $(t_{1},\cdots,t_{n})\in\mathbb{R}^{n}$.
@@
It is not hard to see that $v\in C^{\beta}_{loc}\cap W^{1,q}_{loc}\cap W^{2,\,p}_{loc}\,(q<\frac{2}{1-\beta},\,p<\frac{2}{2-\beta})$, and $w\in C^{\gamma}_{loc}\cap W^{1,q}_{loc}\cap W^{2,p}_{loc}\,(q<\frac{2}{1-\gamma},\,p<\frac{2}{2-\gamma})$ when $0\leq\gamma\leq 1$, $w\in C^{2-\gamma}_{loc}\cap W^{1,q}_{loc}\cap W^{2,p}_{loc}\,(q<\frac{2}{\gamma-1},\,p<\frac{2}{\gamma})$ when $21<\gamma<$ 2. Since these two families of solutions include ones with arbitrarily small Hölder exponents, Examples 1.4 and 1.5 show that there is no interior Hölder estimate for equation (1), even under the toric symmetry and entirety conditions.
Priori to our examples, entire solutions on $\mathbb{C}^{n}$ with different singularity profiles were constructed by Błocki [4] and He [21], respectively. Błocki’s solution is $v(z)=\frac{n}{n-1}\left(1+|z_{1}|^{2}\right)|z^{\prime}|^{2-\frac{2}{n}}$, where $z^{\prime}=(z_{2},\cdots,z_{n})$. It is both a pluripotential and viscosity solution to $\det(\partial\bar{\partial}u)=\left(1+|z_{1}|^{2}\right)^{n-2}$ (strong solution for $n\geq 3$). This example, together with Błocki-Dinew’s work [5], jointly show that $W_{loc}^{2,n(n-1)}$ is the critical Sobolev regularity needed to guarantee the higher regularity of the solutions. He [21] provided the example $w(z)=n^{\frac{2}{n}}|z_{1}|^{\frac{2}{n}}\left(1+|z^{\prime}|^{2}\right)$, which also solves equation (1) both in the viscosity and the pluripotential sense. For dimensions $n\geq 3$, He’s solution is non-Lipschitz, indicating that the interior gradient estimate for the complex Monge-Ampère equation fails. Our examples complement this fact for dimension $2$. Note that Błocki’s solution, He’s solution and our solutions for $\beta=0$ and $\gamma=1$ essentially coincide in complex dimension $2$.
@@
Examples 1.2, 1.4 and 1.5 show that there are “continuous” singularity profiles lying between the critical regularity for self-improvements to smoothness and the almost lowest regularity for defining the complex Monge-Ampère operator. Similar “dense” singularity profiles exist for the real Monge-Ampère operator (cf. [7]).
Compared to the singular solutions of the real Monge-Ampère equation, the much more singular profiles of our solutions highlight the elusiveness of the complex Monge-Ampère operator. A notable characteristic of the singular solutions by Błocki, He and ourselves is their explicit nature, which may be attributed to the “double divergence structure” of the complex Monge-Ampère operator in complex dimension two.
@@ 3. Counterexamples to the interior Dini continuity
on $\mathbb{D}_{1-\varepsilon}\times\mathbb{C}$, where $\mathbb{D}_{1-\varepsilon}$ denotes the disc of radius $1-\varepsilon$ centered at the origin in $\mathbb{C}$. For any compact subset $\Omega$ of $\mathbb{D}\times\mathbb{C}$, $u^{\varepsilon}$ converges decreasingly and uniformly to $u$ on $\Omega$ as $\varepsilon\to 0$. Next we show that $u^{\varepsilon}$ is plurisubharmonic and $\det\left(\partial\bar{\partial}u^{\varepsilon}\right)$ converges to $1$ in $L^{1}(\Omega)$ as $\varepsilon\rightarrow 0$. For simplicity we use the notation
@@ 4. Two families of solutions with continuous Hölder exponents
Acknowledgement. The authors would like to thank Weiming Shen, Yu Yuan, Bin Zhou, and Xiaohua Zhu for helpful discussions.

### v2

@@ Abstract.
We present an explicit pluripotential and viscosity solution to the complex Monge-Ampère equation with constant right-hand side on $\mathbb{D}\times\mathbb{C}^{n-1}\,(n\geq 2)$, which lies merely in $W^{1,2}_{loc}\cap W^{2,1}_{loc}$ and is not even Dini continuous. Additionally, we exhibit two families of explicit entire toric solutions on $\mathbb{C}^{n}$ with continuous Hölder exponent $\alpha\in(0,1)$ and $W^{2,p}_{loc}$ exponent $p\in(1,2)$.
@@
Jiaxiang Wang is partially supported by the Fundamental Research Funds for the Central Universities Nankai University (050-63241427). Wenlong Wang is partially supported by NSFC 12071338, 12001292, and the Fundamental Research Funds for the Central Universities Nankai University (050-63243065).
@@ 1. Introduction
Complex Monge-Ampère equations play an important role in complex analysis and geometry. There are two fundamental methods for generalizing the notion of solution to functions with low regularity. One is the pluripotential approach founded by Bedford and Taylor [2], while the other is the viscosity approach introduced by Crandall and Lions for nonlinear PDEs [13] (see the well-known survey [12]) and developed for the complex Monge-Ampère equation in [15, 18, 19, 20, 28].
@@
where ¯$\partial\bar{\partial}u$ denotes the complex Hessian of $u$. It remains an open question whether a generalized solution possesses interior Hölder regularity. For two-dimensional solutions, even the validity of the interior gradient estimate is unknown.
@@
Clearly, Hölder continuity implies Dini continuity. Therefore, Theorem 1.1 indicates that pure interior Hölder estimate fails for the complex Monge-Ampère equation. This theorem also shows that the equation provides almost no regularity improvement for the gradient of the solution without additional regularity assumptions, as a bounded plurisubharmonic function naturally lies in $W^{1,2}_{loc}$. For the Hessian, it remains unclear whether $L^{1}_{loc}$ is the minimal regularity, as is the case for the real Monge-Ampère equation (cf. [14, 25, 26]).
@@
Besides Example 1.2, we also construct examples that are toric and global on ${\mathbb{C}}^{n}$, i.e., functions satisfying $u(z_{1},\cdots,z_{n})=u(e^{{\sqrt{-1}}t_{1}}z_{1},\cdots,e^{{\sqrt{-1}}t_{n}}z_{n})$ for any $(z_{1},\cdots,z_{n})\in\mathbb{C}^{n}$ and $(t_{1},\cdots,t_{n})\in\mathbb{R}^{n}$.
@@
It is straightforward to see that $v\in C^{\beta}_{loc}\cap W^{1,q}_{loc}\cap W^{2,\,p}_{loc}\,(q<\frac{2}{1-\beta},\,p<\frac{2}{2-\beta})$. Similarly, when $0\leq\gamma\leq 1$, $w\in C^{\gamma}_{loc}\cap W^{1,q}_{loc}\cap W^{2,p}_{loc}\,(q<\frac{2}{1-\gamma},\,p<\frac{2}{2-\gamma})$, and when $21<\gamma<$ 2, $w\in C^{2-\gamma}_{loc}\cap W^{1,q}_{loc}\cap W^{2,p}_{loc}\,(q<\frac{2}{\gamma-1},\,p<\frac{2}{\gamma})$. Since these two families of solutions include ones with arbitrarily small Hölder exponents, Examples 1.4 and 1.5 show that there is no interior Hölder estimate for equation (1), even under the toric symmetry and entirety conditions.
Before our examples, entire solutions on $\mathbb{C}^{n}$ with different singularity profiles were constructed by Błocki [4] and He [21], respectively. Błocki’s solution is $v(z)=\frac{n}{n-1}\left(1+|z_{1}|^{2}\right)|z^{\prime}|^{2-\frac{2}{n}}$, where $z^{\prime}=(z_{2},\cdots,z_{n})$. It is both a pluripotential and viscosity solution to $\det(\partial\bar{\partial}u)=\left(1+|z_{1}|^{2}\right)^{n-2}$ (strong solution for $n\geq 3$). This example, together with Błocki-Dinew’s work [5], jointly show that $W_{loc}^{2,n(n-1)}$ is the critical Sobolev regularity needed to guarantee the higher regularity of the solutions. He [21] provided the example $w(z)=n^{\frac{2}{n}}|z_{1}|^{\frac{2}{n}}\left(1+|z^{\prime}|^{2}\right)$, which also solves equation (1) both in the viscosity and the pluripotential sense. For dimensions $n\geq 3$, He’s solution is non-Lipschitz, indicating that the interior gradient estimate for the complex Monge-Ampère equation fails. Our examples complement this fact for dimension $2$. Note that Błocki’s solution, He’s solution and our solutions for $\beta=0$ and $\gamma=1$ essentially coincide in complex dimension $2$.
@@
Examples 1.2, 1.4 and 1.5 show that there are “continuous” singularity profiles lying between the critical regularity needed for self-improvement to smoothness and the near-minimal regularity required for defining the complex Monge-Ampère operator. Similar “dense” singularity profiles exist for the real Monge-Ampère operator (cf. [7]).
Compared to the singular solutions of the real Monge-Ampère equation, the much more singular profiles of our solutions underscore the elusiveness of the complex Monge-Ampère operator. A notable characteristic of the singular solutions by Błocki, He, and ourselves is their explicit form, which may be attributed to the “double divergence structure” of the complex Monge-Ampère operator in complex dimension two.
@@ 3. Counterexamples to the interior Dini continuity
on $\mathbb{D}_{1-\varepsilon}\times\mathbb{C}$, where $\mathbb{D}_{1-\varepsilon}$ denotes the disc of radius $1-\varepsilon$ centered at the origin in $\mathbb{C}$. For any compact subset $\Omega$ of $\mathbb{D}\times\mathbb{C}$, $u^{\varepsilon}$ converges decreasingly and uniformly to $u$ on $\Omega$ as $\varepsilon\to 0$. Next we show that $u^{\varepsilon}$ is plurisubharmonic and $\det\left(\partial\bar{\partial}u^{\varepsilon}\right)$ converges to $1$ in $L^{1}(\Omega)$ as $\varepsilon\rightarrow 0$. For simplicity, we use the notation
@@ 4. Two families of solutions with continuous Hölder exponents
Acknowledgement. The authors would like to thank Weiming Shen, Yuguang Shi, Xujia Wang, Yu Yuan, Bin Zhou, and Xiaohua Zhu for helpful discussions.